# OutLines: Modeling Spectral Lines from Winds, Bubbles, and Outflows

### v1

@@ 2.5 Outflow Properties
Following Flury2023, the characteristic outflow radius $x_{out}$ corresponds to the maximum momentum density $\dot{\rho}_{p}=\mu m_{\rm H}n_{0}v_{\infty}n[x]w[x]$, i.e., where the wind effects are strongest. In the case of pulse-like density distributions, the center of the gas distribution $x_{1}$ will be close to, but not always the same as, $x_{out}$.The radius $x_{out}$ of maximum momentum density scales all other outflow properties, first and foremost the characteristic outflow velocity $v_{out}$ by determining the velocity at $x_{out}$ using the velocity field $w$ such that
$v_{out}=v_{\infty}w[x_{out}].$ (32)
The effective column density $N_{out}$ at the characteristic radius can be determined by integrating over the density profile from the base of the outflow ($x=1$) to the peak momentum density $x=x_{out}$ (e.g., Flury2023, their Equation 13 for the case of a CAK velocity field and power law density profile). Note that $N_{out}$ is not to be confused with $N$ from the absorption line profile, which is the total column density along the line of sight. $N_{out}$ necessarily depends on the space density $n_{0}$ and the radius at the base of the outflow $R_{0}$, which are unconstrained parameters in OutLines. As such, OutLines computes the relative column density $\mathcal{R}$ where
$\frac{N_{out}}{n_{0}R_{0}}=\mathcal{R}=\int\limits_{1}^{x_{out}}n[x]{\rm d}x.$ (33)
@@
The mass outflow rate $\dot{M}$ is given by the momentum flux $\mu m_{\rm H}n_{0}v_{out}$ with a geometric projection $4\pi R_{0}^{2}(\cos\theta_{c}-\cos\theta_{o})$ such that
$\dot{M}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}x_{out}\mathcal{R}\mu m_{\rm H}v_{out},$ (34)
a more general case of Equation 12 from Flury2023. Such formalism is frequent in the literature with differences in implementation of the geometric projection (e.g. 2005ApJS..160...87R; 2005ApJS..160..115R; Crenshaw2012; Maiolino2012; Fiore2017; Luminari2018; Luminari2024A&A; Flury2023). The outflow momentum rate $\dot{p}$ expresses the rate at which momentum is injected into the outflow. Typically, $\dot{p}$ is defined as the product of outflow velocity and mass outflow rate (e.g., 2005ApJS..160..115R, their Equation 9) such that
$\dot{p}_{out}=v_{out}\dot{M}_{out}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}v_{out}^{2}.$ (35)
The outflow mechanical luminosity $\dot{E}$ expresses the injection of kinetic energy into the outflow. Typically, $\dot{E}$ is defined as the product of outflow velocity and mass outflow rate (e.g., 2005ApJS..160..115R, their Equation 11) such that
$\dot{E}_{out}\begin{split}=&\frac{1}{2}\dot{M}\left(v_{out}^{2}+3\sigma_{v}^{2}\right)\\ =&2\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}\left(v_{out}^{3}+3v_{out}\sigma_{v}^{2}\right).\end{split}$ (36)
These values can be readily computed using constraints on $x_{out}$ and $v_{\infty}$ and possible geometric terms $\theta_{o}$ and $\theta_{c}$, all obtained from the line profile. The only two unknowns are $n_{0}$ and $R_{0}$, which cannot be directly informed by the line profile and require additional measurements (spatial for $R_{0}$, spectral for $n_{0}$) or assumptions based on prior knowledge. For instance, $n_{0}=10\rm~cm^{-3}$ and $R_{0}=0.5$ kpc may be reasonable for galactic outflows (e.g., 1985Natur.317...44C; Flury2023).
@@
$\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=0.191\mathcal{R}[\cos\theta_{c}-\cos\theta_{o}]\left[\frac{v_{out}}{\rm km~s^{-1}}\right]$ (37)
$\frac{\dot{p}}{\rm dyne}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-3.2}\left[\frac{v_{out}}{\rm km~s^{-1}}\right]\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (38)
$\frac{\dot{E}}{\rm erg~s^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-6.5}\left[\frac{v_{out}}{\rm km~s^{-1}}\right]^{2}\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (39)
Typically, $n_{0}R_{0}^{2}$ is of order unity for the case studies in §4, with values ranging from about 1 to 5 cm-3 kpc2, although for Mrk 462, Flury2023 find $n_{0}R_{0}^{2}=12$ cm-3 kpc2.
@@ 3 The OutLines Code
OutLines, a portmanteau of outflows and spectral lines, implements all of the above model components in a single code for emission and absorption lines. The code is designed to be simultaneously efficient and user-friendly, with a straightforward user interface and ready scientific application. Below, we assess OutLines in the contexts of standard techniques, including empirical percentiles and multi-Gaussian fitting. Details of the architecture and use of the OutLines code can be found in Appendix C.
@@ 3.1 Empirical Percentiles
Figure 14: Velocity percentiles (black dashed lines) for an arbitrary emission (top) and absorption (bottom) line (purple) generated by OutLines for a directed, filled cone geometry (Geometry=’FilledCones’), default parameters, and otherwise default settings. The cumulative distribution function (fraction of integrated flux) is shown in red. For reference, we show the characteristic outflow velocity $v_{out}$ (Equation 32) in blue.
@@ 3.2 Modeling Line Profiles
To facilitate modeling of observed line profiles, the Profile classes contain several attributes and functions. Users can access the model line profile function directly via the Profile attribute, which receives rest wavelengths and parameters from the user and returns the total line profile predicted by the OutLines code. When instantiated, the Profile classes also contain initial parameter guesses and boundaries accessible via the get_params() and get_bounds() methods. For convenience, these two classes also contain the init_params function which will sample walkers within the current parameter bounds given the current parameter values. These initial parameter guesses can be readily seeded into an MCMC sampler while the boundaries can be folded into a prior. For this purpose, the log_probability() bound method includes both the likelihood and a uniform bounded prior incorporating the boundary values contained in the object. Users are able to update any or all of these bound methods as needed. An example of modeling with OutLines is provided in Appendix E.
@@
With this caveat in mind, we compare the results between the standard method and our new approach with OutLines in modeling an outflow. OutLines is able to recover all input parameters within the imposed 10% uncertainty due to noise. The double Gaussian model over-predicts the outflow flux by a factor of two and under-predicts the static gas flux by a factor of three while characterizing the outflow as red-shifted relative to the systemic velocity of the static gas. The double Gaussian model results lead to vastly different physical interpretations of the observations than those from OutLines: one would conclude the presence of a more massive outflow that is launched from a location traveling away from the static gas relative to the observer at $100$ $\rm km~s^{-1}$ but at characteristic velocities of $v_{out}=100$ $\rm km~s^{-1}$. One might even be inclined to interpret the double-Gaussian result as evidence for an inflow rather than an outflow, which is not the case. The OutLines best-fit model indicates an outflow launched from the same systemic velocity as the static gas with a bidirectional cone geometry and with characteristic velocities of $v_{out}=220$ $\rm km~s^{-1}$. Following Equation 34, accounting for the geometry (a factor of $1/(1-\cos\pi/6)\approx 7.5$), differences in line flux ($n_{0}\propto\sqrt{F_{0}}$ gives a factor $\sqrt{2}$), and difference in $v_{out}$ (factor of 0.5), the double Gaussian model yields a mass loss rate higher than that of OutLines by a factor of five. Thus, while the velocity inferred from the double Gaussian model is half that from OutLines, the double Gaussian model implies a substantially higher impact of the outflow on its environment than the OutLines model.
@@ 3.3 Absorption vs Emission
A key result of the development and applications of OutLines is that absorption and emission lines optimize constraints on different properties of outflows. Absorption lines are more sensitive to the line-of-sight, as evidenced in the line profiles (Figures 2 and 4) and the narrow range in velocities contributing to each wavelength bin due to the “down-the-barrel” effect (Figures 7 and 9), making them good probes of the density profile and velocity field but poor tracers of the geometry. Emission lines trace the global gas distribution since velocities from across the majority of the gas contribute to each velocity bin (Figure 9), making them good probes of the geometry (Figures 11, 12, 13) but, as a result, lose some constraining power on the density profile and velocity field. As such, combined modeling of both emission and absorption provides optimal constraints on the outflow properties.
@@ 4 A High Capacity for Science
All modeling with OutLines is done here using emcee to sample the posterior on the parameters. The MCMC chain consists of a number of walkers equal to twice the number of free parameters plus one (the OutLines default), with each walker taking $10^{3}$ steps with an additional $200$ burn-in steps given the number of walkers (minimum of 10 walkers for at least 2000 individual burn-in steps), which we find to be sufficient to “forget” the initial guess parameters. The initial guess parameters, unless otherwise stated, are obtained using the scipy.optimize.least_squares trust-reflective algorithm fit. The prior is assumed to be uniform bounded unless otherwise stated. We show posterior sampling for each case in Appendix F.
Table 1: Best-fit parameters and properties of test-case objects: Knot A of giant extragalactic H ii region NGC 5471, super star cluster in the center of Green Pea J1044+3053, star-forming galaxy Mrk 1486, and Seyfert 2 NGC 2992. As discussed in §2.5, reported properties of $\dot{M}$, $\dot{p}$, and $\dot{E}$ are normalized to $n_{0}R_{0}^{2}$ with $n_{0}$ in units of cm-3 and $R_{0}$ in kpc (see Equations 37-39).
@@
OutLines model BetaCAK, PowerLaw, Sphere BetaCAK, PowerLaw, Sphere BetaCAK, DampedPulses, Sphere AccPlaw, PowerLaw, HollowCones, Disk
@@
$\alpha$ $0.98$  $\pm$ $0.31$ $1.39$  $\pm$ $0.76$ $1.42$  $\pm$ $0.37$
$\beta$ $4.07$  $\pm$ $1.24$ $2.27$  $\pm$ $1.10$ $0.71$  $\pm$ $0.25$ $2.48$  $\pm$ $0.42$
$v_{\infty}$ $\rm km~s^{-1}$] $396$  $\pm$ $75$ $213$  $\pm$ $26$ $624$  $\pm$ $38$ $1034$  $\pm$ $108$
$b$ or $\sigma_{v}$ $\rm km~s^{-1}$] $28$  $\pm$ 11 $34$  $\pm$ 2020 $50$  $\pm$ 2020 $77$  $\pm$ 11
@@
$W_{50}$ (“FWHF”) $\rm km~s^{-1}$] $49$  $\pm$ 11 $87$  $\pm$ 44 $181$  $\pm$ $27$ $127$  $\pm$ 44
$W_{80}$ $\rm km~s^{-1}$] $109$  $\pm$ 33 $147$  $\pm$ 66 $305$  $\pm$ $77$ $289$  $\pm$ $12$
$x_{out}$ $5.16$  $\pm$ $2.04$ $2.63$  $\pm$ $^{1.78}_{1.09}$ $1.24$  $\pm$ $0.17$ $1.35$  $\pm$ $0.02$
$v_{out}$ $165$  $\pm$ $15$ $72$  $\pm$ $17$ $188$  $\pm$ $78$ $597$  $\pm$ $267$
$\dot{M}/n_{0}R_{0}^{2}$ $\rm M_{\odot}~yr^{-1}$] $105$  $\pm$ $59$ $22.1$  $\pm$ $^{33.6}_{14.2}$ $9.8$  $\pm$ $3.2$ $9.5$  $\pm$ $0.4$
$\dot{p}/n_{0}R_{0}^{2}$ $\rm 10^{34}~dyne$] $10.9$  $\pm$ $7.2$ $1.00$  $\pm$ $^{2.21}_{0.71}$ $0.95$  $\pm$ $0.54$ $3.59$  $\pm$ $0.25$
$\dot{E}/n_{0}R_{0}^{2}$ $\rm 10^{42}~erg~s^{-1}$] $0.90$  $\pm$ $0.69$ $0.04$  $\pm$ $^{0.11}_{0.03}$ $0.09$  $\pm$ $0.07$ $1.07$  $\pm$ $0.11$

### v2

@@ 2.5 Outflow Properties
Following Flury et al. (2023, see their Appendix B), the characteristic outflow radius $x_{\rm out}$ corresponds to the maximum momentum density $\dot{\rho}_{p}=\mu m_{\rm H}n_{0}v_{\infty}n[x]w[x]$, i.e., where the wind effects are strongest. In the case of pulse-like density distributions, the center of the gas distribution $x_{1}$ will be close to, but not always the same as, $x_{\rm out}$.The radius $x_{\rm out}$ of maximum momentum density scales all other outflow properties, first and foremost the characteristic outflow velocity $v_{\rm out}$ by determining the velocity at $x_{\rm out}$ using the velocity field $w$ such that
$v_{\rm out}=v_{\infty}w[x_{\rm out}].$ (36)
The effective column density $N_{\rm out}$ at the characteristic radius can be determined by integrating over the density profile from the base of the outflow ($x=1$) to the peak momentum density $x=x_{\rm out}$ (e.g., Flury et al., 2023, their Equation 13 for the case of a CAK velocity field and power law density profile). Note that $N_{\rm out}$ is not to be confused with $N$ from the absorption line profile, which is the total column density along the line of sight. $N_{\rm out}$ necessarily depends on the space density $n_{0}$ and the radius at the base of the outflow $R_{0}$, which are unconstrained parameters in OutLines. As such, OutLines computes the relative column density $\mathcal{R}$ where
$\frac{N_{\rm out}}{n_{0}R_{0}}=\mathcal{R}=\int\limits_{1}^{x_{\rm out}}n[x]{\rm d}x.$ (37)
@@
The mass outflow rate $\dot{M}$ is given by the momentum flux $\mu m_{\rm H}n_{0}v_{\rm out}$ with a geometric projection $4\pi R_{0}^{2}(\cos\theta_{c}-\cos\theta_{o})$ such that
$\dot{M}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}v_{\rm out},$ (38)
a more general case of Equation 12 from Flury et al. (2023). Such formalism is frequent in the literature with differences in implementation of the geometric projection (e.g. Rupke et al., 2005a, b; Crenshaw and Kraemer, 2012; Maiolino et al., 2012; Fiore et al., 2017; Luminari et al., 2018, 2024; Flury et al., 2023). The outflow momentum rate $\dot{p}$ expresses the rate at which momentum is injected into the outflow. Typically, $\dot{p}$ is defined as the product of outflow velocity and mass outflow rate (e.g., Rupke et al., 2005b, their Equation 9) such that
$\dot{p}=v_{\rm out}\dot{M}=4\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}v_{\rm out}^{2}.$ (39)
The outflow mechanical luminosity $\dot{E}$ expresses the injection of kinetic energy into the outflow. Typically, $\dot{E}$ is defined as the product of outflow velocity and mass outflow rate (e.g., Rupke et al., 2005b, their Equation 11) such that
$\dot{E}_{out}\begin{split}=&\frac{1}{2}\dot{M}\left(v_{\rm out}^{2}+3\sigma_{v}^{2}\right)\\ =&2\pi(\cos\theta_{c}-\cos\theta_{o})n_{0}R_{0}^{2}\mathcal{R}\mu m_{\rm H}\left(v_{\rm out}^{3}+3v_{\rm out}\sigma_{v}^{2}\right).\end{split}$ (40)
These values can be readily computed using constraints on $x_{\rm out}$ and $v_{\infty}$ and possible geometric terms $\theta_{o}$ and $\theta_{c}$, all obtained from the line profile. The only two unknowns are $n_{0}$ and $R_{0}$, which cannot be directly informed by the line profile and require additional measurements (spatial for $R_{0}$, spectral for $n_{0}$) or assumptions based on prior knowledge. For instance, $n_{0}=10\rm~cm^{-3}$ and $R_{0}=0.5$ kpc may be reasonable for galactic outflows (e.g., Chevalier and Clegg 1985; Flury et al. 2023).
@@
$\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=0.191\mathcal{R}[\cos\theta_{c}-\cos\theta_{o}]\left[\frac{v_{\rm out}}{\rm km~s^{-1}}\right]$ (41)
$\frac{\dot{p}}{\rm dyne}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-3.2}\left[\frac{v_{\rm out}}{\rm km~s^{-1}}\right]\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (42)
$\frac{\dot{E}}{\rm erg~s^{-1}}\left[\frac{n_{0}R_{0}^{2}}{\rm cm^{-3}~kpc^{2}}\right]^{-1}$ $=10^{-6.5}\left[\frac{v_{\rm out}}{\rm km~s^{-1}}\right]^{2}\left[\frac{\dot{M}}{\rm M_{\odot}~yr^{-1}}\right]$ (43)
Typically, $n_{0}R_{0}^{2}$ is of order unity for the case studies in §4, with values ranging from about 1 to 5 cm-3 kpc2, although for Mrk 462, Flury et al. (2023) find $n_{0}R_{0}^{2}=12$ cm-3 kpc2.
@@ 3 The OutLines Code
OutLines, a portmanteau of outflows and spectral lines, implements all of the above model components in a single code for emission and absorption lines. The code is designed to be simultaneously efficient and user-friendly, with a straightforward user interface and ready scientific application. Below, we assess OutLines in the contexts of standard techniques, including empirical percentiles and multi-Gaussian fitting. Details of the architecture and use of the OutLines code can be found in Appendix B with additional examples available in Jupyter notebooks online at https://github.com/sflury/OutLines.
@@ 3.1 Empirical Percentiles
Figure 15: Velocity percentiles (black dashed lines) for an arbitrary emission (top) and absorption (bottom) line (purple) generated by OutLines for a directed, filled cone geometry (Geometry=’FilledCones’) for $i=45^{\circ}$ with $v_{0}=0.1v_{\infty}$ (FromRest=False), default parameters, and otherwise default settings. The cumulative distribution function (fraction of integrated flux) is shown in red. For reference, we show the characteristic outflow velocity $v_{\rm out}$ (Equation 36) in blue.
@@ 3.2 Modeling Line Profiles
To facilitate modeling of observed line profiles, the Profile classes contain several attributes and functions. Users can access the model line profile function directly via the Profile attribute, which receives rest wavelengths and parameters from the user and returns the total line profile predicted by the OutLines code. When instantiated, the Profile classes also contain initial parameter guesses and boundaries accessible via the get_params() and get_bounds() methods. For convenience, these two classes also contain the init_params function which will sample walkers within the current parameter bounds given the current parameter values. These initial parameter guesses can be readily seeded into an MCMC sampler while the boundaries can be folded into a prior. For this purpose, the log_probability() bound method includes both the likelihood and a uniform bounded prior incorporating the boundary values contained in the object. Users are able to update any or all of these bound methods as needed. An example of modeling with OutLines is provided in Appendix D.
@@
With this caveat in mind, we compare the results between the standard method and our new approach with OutLines in modeling an outflow. OutLines is able to recover all input parameters within the imposed 10% uncertainty due to noise. The double Gaussian model over-predicts the outflow flux by a factor of two and under-predicts the static gas flux by a factor of three while characterizing the outflow as red-shifted relative to the systemic velocity of the static gas. The double Gaussian model results lead to vastly different physical interpretations of the observations than those from OutLines: one would conclude the presence of a more massive outflow that is systematically traveling away from the static gas at $15$ $\rm km~s^{-1}$, perhaps merger-like activity. One might even be inclined to interpret the double-Gaussian result as evidence for an inflow rather than an outflow, which is not the case. The OutLines best-fit model indicates an outflow launched from the same systemic velocity as the static gas with a bidirectional cone geometry and with characteristic velocities of $v_{\rm out}=220$ $\rm km~s^{-1}$. Following Equation 38, accounting for the geometry (a factor of $1/(1-\cos\pi/6)\approx 7.5$), differences in line flux ($n_{0}\propto\sqrt{F_{0}}$ gives a factor $\sqrt{2}$), and difference in $v_{\rm out}$ (factor of 0.5), the double Gaussian model yields a mass loss rate higher than that of OutLines by a factor of five. Thus, while the velocity inferred from the double Gaussian model is half that from OutLines, the double Gaussian model implies a substantially higher impact of the outflow on its environment than the OutLines model.
@@ 3.3 Absorption vs Emission
A key result of the development and applications of OutLines is that absorption and emission lines optimize constraints on different properties of outflows. Absorption lines are more sensitive to the line-of-sight, as evidenced in the line profiles (Figures 2 and 4) and the narrow range in velocities contributing to each wavelength bin due to the “down-the-barrel” effect (Figures 7 and 9), making them good probes of the density profile and velocity field but poor tracers of the geometry. Emission lines trace the global gas distribution since velocities from across the majority of the gas contribute to each velocity bin (Figure 9), making them good probes of the geometry (Figures 11, 12, 14) but, as a result, lose some constraining power on the density profile and velocity field. As such, combined modeling of both emission and absorption provides optimal constraints on the outflow properties.
@@ 4 A High Capacity for Science
All modeling with OutLines is done here using emcee to sample the posterior on the parameters. The MCMC chain consists of a number of walkers equal to twice the number of free parameters plus one (the OutLines default), with each walker taking $10^{3}$ steps with an additional $200$ burn-in steps given the number of walkers (minimum of 10 walkers for at least 2000 individual burn-in steps), which we find to be sufficient to “forget” the initial guess parameters. The initial guess parameters, unless otherwise stated, are obtained using the scipy.optimize.least_squares trust-reflective algorithm fit. The prior is assumed to be uniform bounded unless otherwise stated. We show posterior sampling for each case in Appendix E.
Table 1: Best-fit parameters and properties of test-case objects: Knot A of giant extragalactic H ii region NGC 5471, super star cluster in the center of Green Pea J1044+3053, star-forming galaxy Mrk 1486, and Seyfert 2 NGC 2992. As discussed in §2.5, reported properties of $\dot{M}$, $\dot{p}$, and $\dot{E}$ are normalized to $n_{0}R_{0}^{2}$ with $n_{0}$ in units of cm-3 and $R_{0}$ in kpc (see Equations 41-43).
@@
OutLines model BetaCAK, PowerLaw, Sphere BetaCAK, PowerLaw, Sphere, Launched BetaCAK, DampedPulses, Sphere AccPlaw, PowerLaw, HollowCones, Disk
@@
$\alpha$ $0.23$  $\pm$ $0.05$ $0.58$  $\pm$ $0.50$ $0.03$  $\pm$ $0.06$
$\beta$ 77  $\pm$ 2 $1.87$  $\pm$ $1.09$ $0.75$  $\pm$ $0.27$ $2.02$  $\pm$ $0.17$
$v_{\infty}$ $\rm km~s^{-1}$] $309$  $\pm$ $14$ $195$  $\pm$ $17$ $567$  $\pm$ $42$ $925$  $\pm$ $65$
$b$ or $\sigma_{v}$ $\rm km~s^{-1}$] $30$  $\pm$ $0.5$ $5.0$  $\pm$ $2.4$ $64$  $\pm$ 2020 $82$  $\pm$ 11
@@
$W_{50}$ (“FWHF”) $\rm km~s^{-1}$] $47$  $\pm$ 11 $88$  $\pm$ 44 $184$  $\pm$ $27$ $235$  $\pm$ $28$
$W_{80}$ $\rm km~s^{-1}$] $105$  $\pm$ 22 $150$  $\pm$ 66 $318$  $\pm$ $77$ $541$  $\pm$ $35$
$x_{\rm out}$ $43$  $\pm$ $23$ $4.22$  $\pm$ $^{11.83}_{~~2.52}$ $1.24$  $\pm$ $0.17$ 2020  $\pm$ $0.02$
$v_{\rm out}$ $246$  $\pm$ 66 $118$  $\pm$ $32$ $188$  $\pm$ $78$ $883$  $\pm$ $46$
$\dot{M}/n_{0}R_{0}^{2}$ $\rm M_{\odot}~yr^{-1}$] $777$  $\pm$ $295$ $44.5$  $\pm$ $^{268.0}_{~~32.7}$ $4.8$  $\pm$ $3.2$ $721$  $\pm$ $321$
$\dot{p}/n_{0}R_{0}^{2}$ $\rm 10^{34}~dyne$] $120$  $\pm$ $47$ $3.21$  $\pm$ $^{27.00}_{~~2.67}$ $0.57$  $\pm$ $0.54$ $430$  $\pm$ $168$
$\dot{E}/n_{0}R_{0}^{2}$ $\rm 10^{42}~erg~s^{-1}$] $15$  $\pm$ 66 $0.19$  $\pm$ $^{0.16}_{2.09}$ $0.05$  $\pm$ $0.05$ $195$  $\pm$ $70$